# Improving stabilizer approximation with quantum strategy

### v1

@@ II The CHSH game
where $\oplus$ is logical XOR, $\wedge$ is logical AND. Assuming the queries are uniformly chosen, we could easily calculate the classical bias and winning probability. To do this, we linearly transform the binary values of $a$ and $b$ into spin values $\{+1,-1\}$, but still use the same labels to denote them. Then, we have for the bias
@@
The CHSH game has a nice interpretation in terms of the stabilizer formalism, as shown in [15]. We slightly modify the notations and presentation in [15] to show the significance for the ground state problem.
Say we have a Hamiltonian:
$H_{0}=-X\otimes X-X\otimes Z+Z\otimes X-Z\otimes Z.$ (8)
To approximate its ground state with stabilizer states, we choose a commuting subset of terms, say $X\otimes X,Z\otimes Z$, as stabilizers. The corresponding stabilizer state is $|\Phi\rangle$, and the corresponding energy, or expectation value, of the Hamiltonian is $-2$. Now we make the following transformation:
@@
$X^{\prime}=\frac{X+Z}{\sqrt{2}}=H,\quad Z^{\prime}=\frac{Z-X}{\sqrt{2}}=\tilde{H}.$ (12)
Then we could rewrite the Hamiltonian as
$H_{0}=-\sqrt{2}(X\otimes X^{\prime}+Z\otimes Z^{\prime}).$ (13)
Now if we choose $\{X\otimes X^{\prime},Z\otimes Z^{\prime}\}$ as our new stabilizers, we get a lower energy $-2\sqrt{2}$. In fact, this is the lowest eigenvalue of $H_{0}$. The corresponding new stabilizer state is simply [15]
$|\Phi^{\prime}\rangle=I\otimes R_{y}^{\dagger}(\frac{\pi}{4})|\Phi\rangle.$ (14)
It is not difficult to check that $X\otimes X^{\prime}$ and $Z\otimes Z^{\prime}$ indeed stabilize $|\Phi^{\prime}\rangle$. Therefore, the quantum strategy used in the CHSH game helps us improve the stabilizer approximation for the ground state energy. The essential point behind the strategy is that, for different qubits we could use completely different sets of Pauli operators to define the stabilizers. In other words, we could gauge the Pauli operators for each qubit individually, with the single-qubit rotations. I am not sure whether “gauge” is the proper word here to describe the procedure exactly, but I prefer to use it anyhow.

### v2

@@ II The CHSH game
where $\oplus$ is logical XOR, $\wedge$ is logical AND. Assuming the queries are uniformly chosen, we could easily calculate the classical bias and winning probability. To do this, we linearly transform the binary values of $a$ and $b$ into parity/spin values $\{+1,-1\}$, but still use the same labels to denote them. Then, we have for the bias
@@
The CHSH game has a nice interpretation in terms of the stabilizer formalism, as shown in [15]. In particular, it is shown that one could encode all the information of the game into a single Hamiltonian, and seek its ground state for the answer. The recipe is as follows. For each question instance $ij$, one includes a Pauli term into the Hamiltonian, with a $X$ factor for the “0” question, and a $Z$ factor for the “1” question. The expected answer, $i\wedge j$, is encoded into the coefficient of the Pauli term. Explicitly, one transforms it into a spin value, and sets the coefficient as its opposite. Doing so, we get the CHSH Hamiltonian [15]:
$H_{0}=-X\otimes X-X\otimes Z-Z\otimes X+Z\otimes Z.$ (8)
Now we try to approximate its ground state with stabilizer states. We could choose a commuting subset of terms, say $X\otimes X,-Z\otimes Z$, as stabilizers. The corresponding stabilizer state is
$|\Psi\rangle=\frac{|01\rangle+|10\rangle}{\sqrt{2}}.$ (9)
And the corresponding energy, or expectation value of the Hamiltonian, is $-2$. The previous quantum strategy inspires us to make the following transformation:
@@
$X^{\prime}=\frac{X+Z}{\sqrt{2}}=H,\quad Z^{\prime}=\frac{Z-X}{\sqrt{2}}=\tilde{H},$ (13)
as in eq.(4). Then we could rewrite the Hamiltonian as
$H_{0}=-\sqrt{2}(X\otimes X^{\prime}-Z\otimes Z^{\prime}).$ (14)
Thus if we choose $\{X\otimes X^{\prime},-Z\otimes Z^{\prime}\}$ as our new stabilizers, we get a lower energy $-2\sqrt{2}$. In fact, this is the lowest eigenvalue of $H_{0}$. The corresponding new stabilizer state is simply [15]
$|\Psi^{\prime}\rangle=I\otimes R_{y}^{\dagger}(\frac{\pi}{4})|\Psi\rangle.$ (15)
It is not difficult to check that $X\otimes X^{\prime}$ and $-Z\otimes Z^{\prime}$ indeed stabilize $|\Psi^{\prime}\rangle$. Therefore, the quantum strategy used in the CHSH game helps us improve the stabilizer approximation for the ground state energy. The essential point behind the strategy is that, for different qubits we could use completely different sets of Pauli operators to define the stabilizers. In other words, we could gauge the Pauli operators for each qubit individually, with the single-qubit rotations. I am not sure whether “gauge” is the proper word here to describe the procedure exactly, but I prefer to use it anyhow.